OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form.

OpenStudy (anonymous):

The formula you'll want to use here is: \[y-y_1=m(x-x_1)\]where m is the slope, and x_1,y_1 are the coordinates of any point on the line. You have the slope, and a point, so you have everything you need to plug into this equation. Then you can simplify if you need.

OpenStudy (anonymous):

y – 5 = 3(x – 2)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@topstryker The point-slope form of the line is y-y1=m(x-x1) Substitute the coordinate values from the given point and the given slope value y-5=3(x-2)

OpenStudy (ybarrap):

you got it @topstryker